OpenStudy (anonymous):

Need Helpp Onn Mathh (: <3

OpenStudy (anonymous):

we are waiting :P

OpenStudy (anonymous):

2x = 20 or x = 10

OpenStudy (anonymous):

Find the equastion in the slope-intercept form the line containing the two point show work. ( -3,13 ) And (-1,-1 )

OpenStudy (anonymous):

slope = (y2 - y1)/(x2 - x1) = (13 - -1)/(-2 - -1) = 14/-1 = -14 so with y = mx + b we have y = -14x + b and using (-1, -1) we have -1 = 14 +b or b = -15 therefore y = -14x - 15